# Efficient tau-pair invariant mass reconstruction with simplified matrix element techniques

### v2

@@ 2.5 Regularization
Following [4], we also consider the Penalized Maximum Likelihood, which is used to reduce the long tails of reconstructed di-tau invariant mass. The motivation for this is the reduction of the long mass tails of reconstructed mass, which can affect the searches for heavy neutral bosons. For this purpose, one could add a penalty term to the likelihood:
$\mathcal{L^{\text{reg}}}(m_{\text{test}}|\text{data})=\mathcal{L}(m_{\text{test}}|\text{data})+r(\Theta)$ (2.20)
@@
  • test mass scaling:
    $m_{\text{test}}\rightarrow\alpha m_{\text{test}}$ (2.21)
@@
Parameter $\alpha$ modifies the peak position of the reconstruction, and $\beta$ modifies both the peak and the width of the reconstructed masses. Both parameters were chosen with a trial and error to give a similar performance as ”cSVfit” and set to $\alpha=1/1.15$, $\beta=6$. However, we acknowledge that most likely this is not the most optimal choice, and the values of $\alpha$ and $\beta$ might be tuned for each decay mode separately.
2.6 Likelihood maximization
Instead of scanning likelihood in the $m_{\tau\tau}$ space, we perform the scan in the ($x_{1},x_{2})$ space, where each point determines $m_{\tau\tau}$. Points with the same value $m_{\tau\tau}$ differ by $\overline{E}_{T}^{\text{true}}$ hypothesis.
We use the brute force grid search method and evaluate the above formula on a parameter grid of $x_{1}$ and $x_{2}$. We choose a grid with a resolution of $100\times 100$. We select the $x_{1}$ and $x_{2}$ values that yield the highest likelihood and use them to compute the momenta of tauons (eq. 2.2) and then the invariant mass of the system:
$m^{2}_{\tau\tau}=\left(\frac{E^{\text{vis}}_{1}}{x_{1}}+\frac{E^{\text{vis}}_{2}}{x_{2}}\right)^{2}-\left(\frac{p^{\text{vis}}_{1}}{x_{1}}+\frac{p^{\text{vis}}_{2}}{x_{2}}\right)^{2}=\frac{m^{2}_{\text{vis}}}{x_{1}x_{2}}$ (2.23)
An example of the likelihood scan is shown in Figure 3. The figure also includes a comparison of two main components of the likelihood: the integral of the phase space and the transfer function.
The brute force scan of the likelihood provides us a full map of the likelihood as a function of the $x_{1}$ and $x_{2}$ parameters. This allows us to estimate the uncertainty of the mass reconstruction in a time-efficient way using the likelihood profile method. We make use of the Wilks [19] theorem, which states that:
@@ 2.7 Uncertainty estimation
The convenience of this method lies in the fact that it depends only on the value of the maximum of the likelihood. Therefore, it can be used without calculating the normalization factor in the likelihood. We use the likelihood maps to find the contour of the region with the masses of the assumed confidence level. An example of contour is shown in Fig. 3. The $x_{1,2}$ values are translated to the mass values using eq. 2.23. The confidence level is spanned by the minimal and maximal values of the mass.
2.8 Additional mass constraints
The way the likelihood is calculated allows for convenient addition of additional constraints. If the analyzed events are expected to originate from a known resonance, such as the $Z^{0}$ or the Higgs boson, one can add a mass constraint to the likelihood to improve the $x_{1,2}$ resolution and therefore the individual $\tau$ energies. The directions can not be constrained, as the algorithm assumes that the taus are collinear with measured visible products. The mass constraint is incorporated by multiplying the likelihood function by a new component, $C_{m}$, which modifies the main likelihood:
$\mathcal{L}(m_{\text{test}}|\text{data})=\mathcal{L}_{b}(m_{\text{test}}|\text{data})\,C(m),$ (2.25)
where $\mathcal{L}_{b}(m_{\text{test}}|d)$ represents the baseline likelihood, as defined in Eq. 2.19.
@@ 2.8 Additional mass constraints
where $\sigma$ is chosen arbitrarily. When selecting $\sigma$, one must balance between a constraint that is too weak, rendering its effect negligible, and one that is too strong, turning the reconstruction algorithm into a pure discriminator. We have evaluated the performance of the algorithm using the transverse momentum ($p_{T}$) of the reconstructed taus as an indicator. This choice is motivated by the fact that information about the $p_{T}$ of individual tau leptons is crucial in $H\rightarrow\tau^{+}\tau^{-}$ analyses, such as CP phase measurements [7].
@@
Figure 4: Time needed to calculate the masses for a given number of cases. The figure shows a comparison between the NumPy implementation (blue line), using batch processing (orange line), and both batch and multiprocessing (green line).
Figure 5: Reconstruction of the $Z^{0}$ (left) and Higgs (right) bosons’ mass with the fastMTT algorithm. The x-axis represents the invariant mass of the reconstructed di-tau system, while the y-axis represents event density.
Figure 6: Relative resolution of $\tau\tau$ reconstruction for $Z^{0}$ (left) and Higgs (right) bosons. The x-axis represents the relative resolution of the mass reconstruction, while the y-axis represents event density.
@@ 3.1 Implementation and time efficiency
Our study shows that using only the fastMTT algorithm, one can process ten thousand events in $46$ seconds, which corresponds to a time budget of $4.6\,\mathrm{ms}$ per event. This is already approximately $50$ times faster than the performance of the cSVfit algorithm.
Utilizing batch processing – necessary for handling a large number of events due to the algorithm’s vectorized implementation – provides an additional speed-up, allowing the same number of events to be processed in approximately $28$ seconds ($2.8\,\mathrm{ms}$ per event). Further acceleration can be achieved by employing multiprocessing with 8 workers, reducing the computation time to 7 seconds, equivalent to $0.7\,\mathrm{ms}$ per event, which is sufficient for the practical needs of physical analysis.
3.2 Mass resolution
@@
The distribution of reconstructed mass is shown in Fig. 5, while the pulls: $(m_{\text{reco}}-m_{\text{true}})/{m_{\text{true}}}$ are shown in Fig. 6.
The mass resolutions, defined as the standard deviation of the reconstructed mass distribution, are $22$ GeV and $35$ GeV for the $Z^{0}$ and Higgs bosons, respectively. These resolutions are comparable to the results obtained in [4].
In order to test whether the algorithm is feasible for searches of new neutral resonances, we have evaluated the algorithm for $\tau\tau$ resonances with masses of $150,175,200,300,500,700$ GeV. We calculated the mean and standard deviation for the testing sample of each mass and plotted them in Figure 7. In this case, one can see that the algorithm is capable of correct reconstruction of masses for low-mass resonances (up to masses of $300$ GeV). In the case of higher masses, the reconstruction of mass is becoming worse; however, it is still sensitive to the mass of the initial neutral boson.
@@
Finally, we have tested the mass uncertainty estimation by analyzing the standardized distribution of the mass:
$z=\frac{x_{\text{reco}}-x_{\text{true}}}{\sigma_{\text{reco}}}.$ (3.1)
The $z$ should asymptotically achieve normal distribution [14]. We use this fact to test the results and compare the pull distribution with the Gaussian function in Fig. 8. To make a comparison, we calculate its Kullback-Leibler divergence [13] from normal distribution with the same mean and covariance $\mathcal{N}(\mu_{z},\sigma_{z})$, obtaining a value $D_{KL}=0.2012$. The $z$ distribution differs from a normal distribution not only in its mean and width, but also in shape. When compared to a Gaussian with the same mean and variance, it appears visibly narrower, suggesting that a simple calibration by scaling the reconstructed uncertainties would lead to their overestimation.
Figure 8: Normalized pull compared to the normal distribution. Normalization was done by subtracting the mean and rescaling by its standard deviation. Their Kullback-Leibler divergence is calculated to be $0.2012$.
@@
Figure 9: Relative resolution of the reconstructed $p_{T}$ without mass constraint (blue) and with constraint given by normal distribution with $\sigma=7.2$ GeV (red).
@@ 4 Summary
A fast version of the matrix element-based algorithm for reconstruction of the di-tau invariant mass has been developed. This algorithm shows similar performance, in terms of mass resolution, with the cSVfit algorithm, but is significantly faster. Access to the full likelihood map in the space of fractions of energy carried by the visible products of the $\tau$ decay allows for easy addition of additional constraints, such as the mass constraint, and event-by-event reconstructed mass uncertainty estimation.

### v3

@@ 2.5 Regularization
Following [10], we also consider a penalized maximum-likelihood approach, introduced to suppress the long tails of the reconstructed di-$\tau$ invariant-mass distribution. The motivation for introducing this penalty is to reduce the extended high-mass tails originating from reconstructed $Z^{0}$ boson events, which would otherwise degrade the sensitivity of searches for heavy neutral resonances. This effect is particularly important because the $Z^{0}$ boson has a much larger production cross section and therefore constitutes a dominant background that must be properly accounted for. To this end, Ref. [10] introduces an additional penalty term in the likelihood:
$\mathcal{L^{\text{reg}}}(\text{data}|x_{1},x_{2})=\mathcal{L}(\text{data}|x_{1},x_{2})+r(\Theta).$ (2.20)
@@
  • test-mass scaling:
    $m_{\text{test}}\rightarrow\alpha m_{\text{test}},$ (2.21)
@@
in the Jacobi factor in equations 2.16 - 2.18.
The parameter $\alpha$ modifies the peak position of the reconstruction, and $\beta$ modifies both the peak and the width of the reconstructed masses.
Both parameters were chosen using a trial-and-error procedure to achieve a performance comparable to that of ”cSVfit”. The parameter $\alpha$ was fixed to $\alpha=1/1.1$, while $\beta$ was chosen depending on the decay channel: $\beta=6$ for fully hadronic decays of two $\tau$ leptons, $\beta=2$ for semi-leptonic decays (one leptonic and one hadronic $\tau$ decay), and $\beta=3.5$ for fully leptonic decays. The comparison is performed using simulated $Z^{0}\to\tau\tau$ events and signal samples of $H\to\tau\tau$, as described in Section 3.2.
We note, however, that the use of constant factors is likely not optimal. The values of $\alpha$ and $\beta$ could be further optimized for specific experimental workflows, in particular for different $\overline{E}_{T}^{\text{miss}}$ reconstruction techniques.
2.6 Likelihood maximization
Instead of scanning the likelihood in the $m_{\tau\tau}$ space, we perform the scan in the ($x_{1},x_{2})$ space, where each point determines $m_{\tau\tau}$ as for Eq. 2.9. Points with the same value $m_{\tau\tau}$ differ by $\overline{E}_{T}^{\text{test}}$ hypothesis according to Eq. 2.14.
We use the grid search method and evaluate the above formula on a parameter grid of $x_{1}$ and $x_{2}$. We use a grid with a resolution of $100\times 100$. We select the $x_{1}$ and $x_{2}$ values that yield the highest likelihood and use them to compute the momenta of taus (eq. 2.2) and then the invariant mass of the system:
$m^{2}_{\text{test}}=\left(\frac{E^{\text{vis}}_{1}}{x_{1}}+\frac{E^{\text{vis}}_{2}}{x_{2}}\right)^{2}-\left(\frac{p^{\text{vis}}_{1}}{x_{1}}+\frac{p^{\text{vis}}_{2}}{x_{2}}\right)^{2}=\frac{m^{2}_{\text{vis}}}{x_{1}x_{2}}.$ (2.23)
An example of the likelihood scan is shown in Figure 3. Fig. 3 contains a comparison of two main components of the likelihood: the integral of the phase space and the transfer function.
@@ 2.6 Likelihood maximization
2.7 Uncertainty estimation
The grid scan of the likelihood provides full map of the likelihood as a function of the $x_{1}$ and $x_{2}$ parameters. This allows to estimate the uncertainty of the mass reconstruction in a time-efficient way using the likelihood profile method. We make use of the Wilks [15] theorem, which implies that:
@@ 2.7 Uncertainty estimation
As the derived confidence levels correspond to the simplified likelihood (following Sec. 2.4), we expect that they may be either over- or underestimated. Therefore, we propose calibrating them using the following procedure.
First, the likelihood is evaluated using simulated Monte Carlo samples of Higgs and $Z^{0}$ boson events. Next, the distribution and variance of the normalized residual (pull) are computed:
$z=\frac{m_{\text{reco}}-m_{\text{true}}}{\sigma_{\text{reco}}},$ (2.25)
where $m_{\text{reco}}$ denotes the reconstructed mass, $m_{\text{true}}$ the true mass, and $\sigma_{\text{reco}}$ the estimated uncertainty of the reconstructed mass.
The quantity $z$ should asymptotically follow a normal distribution [16] $\mathcal{N}(\mu_{z},\sigma_{z})$. Therefore, the reconstructed uncertainty is adjusted according to
$\sigma_{\text{calibrated}}=\sigma_{\text{reco}}\,\sigma_{z}.$ (2.26)
2.8 Additional mass constraint
The likelihood formulation enables the straightforward incorporation of additional constraints, thereby extending its application from background discrimination to energy reconstruction.
Suppose the analyzed events are expected to originate from a known resonance, such as the $Z^{0}$ or the Higgs boson. In that case, one can add a mass constraint to the likelihood to improve the $x_{1,2}$ resolution and therefore the individual $\tau$ energies. The directions cannot be constrained, as the algorithm assumes that the taus are collinear with measured visible products.
The mass constraint is incorporated by multiplying the likelihood function by a new component, $C(m)$, which modifies the main likelihood:
$\mathcal{L}(\text{data}|x_{1},x_{2})=\mathcal{L}_{b}(\text{data}|x_{1},x_{2})\,C(m),$ (2.27)
where $\mathcal{L}_{b}(\text{data}|m_{\text{test}})$ represents the baseline likelihood, as defined in Eq. 2.19 and $m$ is a mass of selected resonance (e.g. $125$ GeV for Higgs).
@@ 2.8 Additional mass constraint
where $\sigma$ is chosen arbitrarily. When selecting $\sigma$, one must balance between a constraint that is too weak, rendering its effect negligible, and one that is too strong, turning the reconstruction algorithm into a pure discriminator. We have evaluated the performance of the algorithm using the transverse momentum ($p_{T}$) of the reconstructed taus as a figure of merit. This choice is motivated by the fact that information about the $p_{T}$ of individual tau leptons is important in $H\rightarrow\tau^{+}\tau^{-}$ analyses, such as CP phase measurements [3].
@@
The mass constraint effectively acts as a Bayesian prior on the di-$\tau$ invariant mass, encoding prior knowledge of the resonance mass. While this improves the precision of the reconstructed $\tau$-lepton energies, it simultaneously biases the reconstructed mass toward the Higgs mass hypothesis. As a result, the likelihood no longer serves as an effective discriminator between signal and background.
Figure 4: Time needed to calculate the masses for a given number of cases. The figure shows a comparison between the NumPy implementation (blue line), using batch processing (orange line), and both batch and multiprocessing (green line). All computations were performed locally on a system equipped with two AMD Opteron 6320 processors (16 logical cores, x86_64 architecture, 2.8 GHz).
Figure 5: Reconstruction of the $Z^{0}$ (left) and Higgs (right) bosons’ mass with the fastMTT algorithm. The x-axis represents the invariant mass of the reconstructed di-tau system, while the y-axis represents event density. For the Higgs boson, the fractions of fully hadronic, semi-leptonic, and fully leptonic events are $61\%$, $34\%$, and $5\%$, respectively, while for the $Z^{0}$ boson they are $65\%$, $30\%$, and $5\%$.
Table 1: Summary of mass reconstruction performance for different decay channels. Mean and standard deviation are reported for each decay channel. Had-had denotes the fully hadronic decay channel, had-lep the semi-leptonic channel, and lep-lep the fully leptonic channel. The parameter $\beta$ represents the regularization strength defined in Eq. 2.22 and is chosen independently for each decay channel.
Process Mass mean [GeV] Mass std. dev. [GeV] $\beta$
Higgs (had-had) $124.5$ GeV $29.7$ GeV 66
Higgs (had-lep) $124.6$ GeV $32.3$ GeV 22
Higgs (lep-lep) $124.7$ GeV $29.1$ GeV $3.5$
$Z^{0}$ (had-had) $99.7$ GeV $23.8$ GeV 66
$Z^{0}$ (had-lep) $98.2$ GeV $28.5$ GeV 22
$Z^{0}$ (lep-lep) $97.0$ GeV $26.7$ GeV $3.5$
Figure 6: Relative resolution of $\tau\tau$ reconstruction for $Z^{0}$ (left) and Higgs (right) bosons. The x-axis represents the relative resolution of the mass reconstruction, while the y-axis represents event density. For the Higgs boson, the fractions of fully hadronic, semi-leptonic, and fully leptonic events are $61\%$, $34\%$, and $5\%$, respectively, while for the $Z^{0}$ boson they are $65\%$, $30\%$, and $5\%$.
Table 2: Summary of the reconstruction resolution, defined as $(m_{\text{reco}}-m_{\text{true}})/m_{\text{true}}$, for different decay channels. Mean and standard deviation are reported for each decay channel. Had-had denotes the fully hadronic decay channel, had-lep the semi-leptonic channel, and lep-lep the fully leptonic channel. The parameter $\beta$ represents the regularization strength defined in Eq. 2.22 and is chosen independently for each decay channel.
Process Resolution mean Resolution std. dev. $\beta$
Higgs (had-had) 0.003-0.003 $0.24$ 66
Higgs (had-lep) 0.003-0.003 $0.26$ 22
Higgs (lep-lep) 0.003-0.003 $0.24$ $3.5$
$Z^{0}$ (had-had) $0.09$ $0.23$ 66
$Z^{0}$ (had-lep) $0.06$ $0.27$ 22
$Z^{0}$ (lep-lep) $0.06$ $0.26$ $3.5$
@@ 3.1 Implementation and time efficiency
Our study shows that using only the fastMTT algorithm, one can process ten thousand events in $46$ seconds, which corresponds to a time budget of $4.6\,\mathrm{ms}$ per event.
Utilizing batch processing – which is necessary for handling a large number of events due to the vectorized implementation of the algorithm – provides an additional speed-up, allowing the same number of events to be processed in approximately $28\,\mathrm{s}$ ($2.8\,\mathrm{ms}$ per event). Further acceleration is achieved by employing multiprocessing with 8 workers, reducing the total computation time to $7\,\mathrm{s}$, corresponding to $0.7\,\mathrm{ms}$ per event, which is sufficient for the practical needs of physics analyses.
The failure rate of the algorithm is negligible.
3.2 Reconstruction resolution
@@
The algorithm was tested in its power to differentiate between signal and background (without mass penalty term). The distribution of reconstructed mass is shown in Fig. 5, while the pulls: $(m_{\text{reco}}-m_{\text{true}})/{m_{\text{true}}}$ are shown in Fig. 6. Results are also reported for different decay channels in Table 1 for reconstructed masses and Table 2 for resolution of the reconstruction.
The mass resolutions, defined as the standard deviations of the reconstructed mass distributions, are $25$ GeV and $30$ GeV for the $Z^{0}$ and Higgs bosons, respectively. The corresponding relative resolutions are $0.25$ for the $Z^{0}$ boson and $0.24$ for the Higgs.
In order to test whether the algorithm is feasible for searches of new neutral resonances, we have evaluated the algorithm for $\tau\tau$ resonances with masses of $150,175,200,300,500,700$ GeV. We calculated the mean and standard deviation for the testing sample of each mass and plotted them in Figure 7. In this case, one can see that the algorithm is capable of correctly reconstructing masses for low-mass resonances (up to masses of $300$ GeV). In the case of higher masses, the reconstruction of mass is becoming worse; however, it still provides an estimator with a linear response to the true resonance mass.
[…]
For fully hadronic decays, we obtained a standard deviation of $1.57$. Since this value exceeds unity, it indicates that the estimated uncertainties systematically underestimate the true mass resolution. In contrast, the semi-leptonic and fully leptonic channels yield standard deviations of $0.93$ and $0.56$ respectively, implying an overestimation of the true uncertainty. Consequently, we calibrate the uncertainty estimates according to Eq. 2.26, treating each decay channel independently.
@@
We examine the calibrated results by comparing the $z$ distribution with a Gaussian function in Fig. 8 for all events and Table 3 for different decay channels. For this comparison, we compute its Kullback–Leibler divergence [22] with respect to a normal distribution $\mathcal{N}(0,1)$, obtaining $D_{KL}=0.121$. The calibrated $z$ distribution differs from a normal distribution not only in its mean but also in its shape. In particular, it appears visibly narrower than $\mathcal{N}(0,1)$, indicating that the estimated uncertainty $\sigma_{\text{reco}}$ is, on average, larger than the actual spread of $m_{\text{reco}}-m_{\text{true}}$ for events close to the true mass (although this may differ between decay channels). This leads to an overestimation of the uncertainty in the vicinity of the true value. At the same time, since the divergence is relatively small, the uncertainty estimate still provides a qualitatively reasonable description of the mass resolution.
Figure 8: Normalized residual compared to the normal distribution, evaluated on the mixed Higgs-$Z^{0}$ sample. Normalization was done by subtracting the mean and rescaling by its standard deviation. Their Kullback-Leibler divergence is calculated to be $0.121$.
Table 3: Summary of the normalized residual (pull) for different decay channels. Mean and standard deviation are reported for each decay channel. Had-had denotes the fully hadronic decay channel, had-lep the semi-leptonic channel, and lep-lep the fully leptonic channel. KL divergence refers to the Kullback–Leibler divergence between the pull distribution and the standard normal distribution $\mathcal{N}(0,1)$.
Process Pull mean Pull std. dev. KL divergence
Higgs (had-had) 0.20-0.20 $1.29$ $0.197$
Higgs (had-lep) 0.17-0.17 $1.06$ $0.141$
Higgs (lep-lep) 0.17-0.17 $0.94$ $0.077$
$Z^{0}$ (had-had) $0.30$ $0.99$ $0.122$
$Z^{0}$ (had-lep) 0.01-0.01 $0.90$ $0.098$
$Z^{0}$ (lep-lep) $0.03$ $0.78$ $0.139$
@@
After applying the mass constraint, the reconstructed $Z^{0}$ mass peaks at $m_{Z^{0}}^{\text{constr}}=117$ GeV with a width of $\sigma_{Z^{0}}=17$ GeV, illustrating the intrinsic bias introduced when the mass observable is replaced by a kinematic constraint to improve the resolution of the $p_{T}$ reconstruction.
Figure 9: Relative resolution of the reconstructed $p_{T}$ of taus without mass constraint (blue) and with constraint given by normal distribution with $\sigma=7.2$ GeV (red). The algorithm was evaluated on the sample coming from the Higgs decay.
3.3 Discussion
To quantify the results, we compare them with [10] using its fastest approach (cSVfit).
In terms of resolution, we obtained a relative reconstruction resolution of 0.230.270.23-0.27 for the $Z^{0}$ and 0.240.260.24-0.26 for the Higgs boson (the ranges reflect results separated into different decay channels). For comparison, cSVfit reports relative resolutions of 0.0810.0810.2640.264 and 0.0790.0790.2680.268, respectively. Given that the present work is based on simplified simulations, we consider these results to be comparable, with only a small loss in resolution in the case of our algorithm.
We observe that our algorithm achieves a comparable performance across different decay channels, whereas cSVfit exhibits improved performance in hadronic channels. This behaviour is expected, since in hadronic decays the $\overline{E}_{T}^{\text{miss}}$ is smaller (as only one neutrino escapes the system), and its impact on the reconstruction is therefore reduced. We attribute the comparable performance observed in our results to the relatively weak $\overline{E}_{T}^{\text{miss}}$ resolution in the simplified simulations used in this study.
For heavy scalar resonances, we observe that our reconstruction remains fairly accurate for masses up to $300$ GeV. For heavier resonances, the reconstructed masses worsen, with a consistent underestimation. This behaviour is likely caused by a combination of poorer $\overline{E}_{T}^{\text{miss}}$ resolution in the high-mass region, limitations of our dataset based on simplified Delphes simulations, and stronger regularization applied to stabilize the results.
When comparing these results with cSVfit, we observe similar behaviour in their algorithm. They report a constant underestimation of the reconstructed $A$ mass, with a mean relative difference of $7.2\%$ in the regularized version of the algorithm (and a mean $2.6\%$ mass overestimation in the unregularized version). We conclude that such underestimation is partially expected and may vary depending on the dataset. One should account for these limitations when performing searches for heavy resonances.
Regarding time efficiency, our algorithm can process one event in approximately $5$ ms in a single-thread evaluation. Compared with cSVfit, which reports processing times of 0.340.340.770.77 s per event, this corresponds to roughly a factor of $100$ speed-up in the case of our algorithm. After comparing the machines used—AMD Opteron 6320 processors (16 logical cores, x86_64 architecture, 2.8 GHz) in our case and a 2.30 GHz Intel® Xeon® E5-2695 v3 for cSVfit – we expect only a small performance difference due to hardware, slightly favouring the machine used in [10]. This represents the main advantage of our approach, which in practical applications may offset the slightly worse resolution performance.
@@ 4 Summary
A fast version of a matrix-element-based algorithm for reconstructing the di-tau invariant mass has been developed. While its mass resolution is slightly worse than that of the cSVfit algorithm, it is significantly faster. Moreover, access to the full likelihood map in the space of energy fractions carried by the visible products of the $\tau$ decay allows for the straightforward inclusion of additional constraints, such as mass constraints, and enables an event-by-event estimation of the reconstructed mass uncertainty.
CRediT authorship contribution statement
Artur Kalinowski: Conceptualization, Methodology, Software, Validation, Writing – review & editing.
Wiktor Matyszkiewicz: Methodology, Software, Validation, Formal analysis, Data curation, Writing – original draft.
Declaration of competing interest(s)
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.